The pdf of a random variable x is given by

f(x)={3/2x2,-1<y<1 both inclusive

{0,elsewhere

Find the pdf of y=(1/2)x2. Hence state the mean and standard deviation of y


1
Expert's answer
2021-07-13T11:48:09-0400

The probability distribution function (PDF) of the given random variable X is

"f_X(x)=\frac{3}{2}x^2", for "-1\leq x\leq 1" and "f_X(x)=0" elsewhere.

The cumulative distribution function (CDF) of X is

"F_X(x)=\int\limits_{-\infty}^{x}f_X(t)dt"

Hence

"F_X(x)=0", for x<1

"F_X(x)=1", for x>1

"F_X(x)=\int\limits_{-1}^x\frac{3}{2}t^2dt=\frac{1}{2}t^3|_{-1}^x=(x^3+1)/2" for "-1\leq x\leq 1".


Consider the random variable "Y=X^2\/2".

"F_Y(y)=P(Y\leq y)=P(X^2/2\leq y)=P(X^2\leq 2y)"

Hence

"F_Y(y)=0", if y<0

"F_Y(y)=1", for 2y>1 (i.e. y>1/2)

"F_Y(y)=F_X(\sqrt{2y})-F_X(-\sqrt{2y})="

"(1+(\sqrt{2y})^3)/2-(1-(\sqrt{2y})^3)/2=2\sqrt{2}y^{3/2}"

Then

"f_Y(y)=\frac{d}{dy}F_Y(y)"

"f_Y(y)=3\sqrt{2}y^{1/2}" for "y\in[0,1/2]" and "f_Y(y)=0" elsewhere.


Expectation of "Y" is "E(Y)=\int\limits_0^{1/2} yf_Y(y)dy=\int\limits_0^{1/2}3\sqrt{2}y^{3/2}dy="

"=3\sqrt{2}\cdot \frac{2}{5}y^{5/2}|_0^{1/2}=3/10=0.3"

Expectation of "Y^2" is "E(Y^2)=\int\limits_0^{1/2} y^2f_Y(y)dy=\int\limits_0^{1/2}3\sqrt{2}y^{5/2}dy="

"=3\sqrt{2}\cdot \frac{2}{7}y^{7/2}|_0^{1/2}=3/28"

Variance of "Y" is "V(Y)=E(Y^2)-E(Y)^2=3\/28-9\/100=3\/175"

Standard deviation is "\sigma_Y=\sqrt{V(Y)}=\sqrt{3/175}=\sqrt{21}/35=0.131"


Answer. "E(Y)=0.3"; "\sigma_Y=0.131".